OpenStudy (anonymous):

Simplify the given expression. (3 + 4i) + (5 − 2i)

OpenStudy (anonymous):

i.e add?

OpenStudy (anonymous):

\(5+3=8,4-2=2\) giving \(8+2i\)

OpenStudy (anonymous):

I got -2+6i but i dont think its correct

OpenStudy (anonymous):

satelite73 is correct. :)

OpenStudy (anonymous):

Oh, I mean satellite73

OpenStudy (anonymous):

Simplify the given expression. (1 + 2i) • (5 − 3i) I also need help with this one

OpenStudy (anonymous):

Using the FOIL method, multiplying the expressions will yield, \[5-3i+10i-6i^{2}\], simplifying, \[7i+11\]

OpenStudy (anonymous):

thats not ananswer choice ..

OpenStudy (dobby1):

@satellite73 is correct

OpenStudy (anonymous):

* an answer

OpenStudy (dobby1):

post the choices

OpenStudy (anonymous):

Its there ! thanks guys ! :] @theANNEbeatable

OpenStudy (dobby1):

np

OpenStudy (anonymous):

No problem! You're welcome! :)